Version 45 (modified by selatham, 14 years ago) (diff)


20 June 2007

  1. Browse status
  2. WMS & WCS
    1. status
    2. Integration(browse)
    3. metadata requirements for testing
    4. decide about next Wed mtg
  3. NDG Closedown timetable

13 June 2007

  1. Security - Critical Path for Beta+Security

All-hands 4th June 2007. Venue Southampton. Time 10:00am to 3:30pm.

  1. Project Status Issues:
    1. Financial and staffing situation
    2. ...and therefore Project Timescale
  2. Technical/Deployment? Status Reports plus potential deliverables
    1. Discovery including NERC and MDIP status (Sue & Bryan )
    2. ...closely followed by Browse (BNL & All)
    3. Vocabulary Services (Roy)
    4. CSML2 (Dominic)
    5. Data Services & Visualisation (Stephen & Dominic)
    6. Security (Phil)
    7. Metadata population - Discovery, Moles and CSML (All)
  3. Lunch
  4. Collaborations requirements (Andrew lead)
    2. NERC Discovery Service
    3. MDIP
  5. Deliverables discussion
    1. inc. indentification of blockers
  6. Decisions/Timetabling?
    1. re-hash tasks etc on trac as go
    2. what do we do in NDG2?
    3. What's left for futures?
  7. AOB

30th May 2007

Note - Sue and Bryan can't make this meeting. If you need them to answer questions then note them for later e-mail or ticketing. Mike away too.

  1. Browse is working with BADC (with simplified xqueries and tuned exist db). Need to get browse working with other DPs ASAP/by Monday. Are there any outstanding obstacles?
    1. PML has working browse, but needs info on tuning exist and is missing some exist libraries ("voclib"?).
    2. BODC has also got working NDG-Browse on a local Sandbox, and copied "xqueryLib" Collection from SVN.
  2. Sample data for CSMLing/overlay. Have all DPs got an idea of what target data they will be CSMLing?
  3. AOB

16th May 2007

9th May 2007

  1. Browse (Aim: to have PML browse working via MOLES basic rather than stub-B)
    • Report (Bryan doesn't think he can make this meeting): As discussed on list, we can't really use MOLES basic either, as that has no names. Kev working on simplified xquery, which we hope we'll have asap. See how that goes.)

2nd May 2007

  1. Report on CSML service meeting last Friday (OwsNotes)
  2. Security schema and MOLES (held over from last week)
  3. Doing away with Stub-B
  4. Browse Status:
    1. Code (need to get rid of stub-b)
    2. Deployment (code:PML, BODC in sandbox, SOC nearly Mod_Python problems, BADC Chinook unstable)
    3. Content (PML best for testing)
  5. Cataloging Coordinate Reference Systems

25th April 2007

  1. Separate security schema from MOLES?
  2. Browse Status
  3. Finance status, end date of current activities.

18th April 2007

  1. DX GUI #647
  2. Eggs and software control
  3. Update on status of security package:
    1. Egg status
    2. DEWS
      • delivery of large datasets
      • Java clients and use of proxy certificates
    3. MyProxy as a  Certificate Authority
    4. NDG readiness: Session Manager, Simple CA, Logging and Login services (PHP)
    5. AHM Paper
  4. (mggr->dom)  CSML raw extract proposal
  5. FastCGI /modpython installations
  6. Status, clarification for ServiceBinding

11th April 2007

No meeting

4th April 2007

  1. Discovery status update.
  2. Objectives for Roy at April 19(30th?) MDIP meeting.
  3. Status, clarification and action plan for implementing ServiceBinding.
  4. FastCGI /modpython installations
  5. NDG security installation status (inc twisted)
  6. CSML status - tooling, creation at !DPs
  7. Planning rapid development of DX/Geosplat replacement. (point to DX/Geoslpat in meantime.)
  8. Date of next meeting (who's away Easter week?)

28th February 2007

  1. Part of the Discovery refactoring is the change across to the official vocab namespaces. This means that all MOLES producers should use these.

14th February 2007

  1. Mod_Python status at SOC?
  2. Completion of SystemStatus page?
  3. Status of MOLES refactoring and ticket destruction?
  4. Status of ServiceBinding?

7th February 2007

  1. Working Grid
    1. Information Status
      1. Where is everyone in their metadata production?
    2. Status and Plans for MOLES
      1. Now that glue has discovery  up, we find that the mini-MOLES conversion have a lot of problems. For examples:
        1. Location keywords become paleo whatsits. (see any NEODC record before and after)
        2. Parameters do not survive the conversion. (see any record).
        3. Actions required: everyone to look at docs and advise Kev on specific problems (this week). Kev to refactor xqueries to support easier adjustment (this week). Kev to fix as many as possible by next week.
    3. Status and plans for discovery
      1. Need to get service binding sorted, ensure MOLES can support it, and ensure that the transformations keep it, then, and only then, will work on browse.
      2. Actions: Bryan to provide service binding final decision (this week).
      3. Status of FastCGI at data providers?
  2. Documentation:
    1. Status of the two-pager. Bryan has provided a revised version to Helen. What next?

Meeting Outcomes

  1. Bryan to finalise ServiceBinding (this week)
  2. Bryan to produce a table of metadata and system status for updating by all parties (DONE)
  3. Bryan to investigate wsgi and mod_python to avoid fastcgi problems at soc (and probably elsewhere). Other options include proxying to lightpd or something else. (DONE, mod_python should be ok, see comment on ticket:639)
  4. Kevin to complete refactoring of xqueries into exist libraries (this week)
  5. All to provide examples of problems with lossy conversions into or out of MOLES (we expect lossy conversion out of MOLES where MOLES has more than the output format, but where the MOLES content ought to end up in the output, we need to tell Kev about it). All problems to be raised as specific tickets. (this week and ongoing)
  6. Sue to get MDIP examples into the discovery exist (this week)
  7. Kev to fix as many of the ticketed problems as possible next week, prioritising any MDIP ones (by next meeting)
  8. Helen to ring Bryan re QUEST and the two-pager (asap).
  9. Sue and Bryan to further investigate memory problems in the discovery ingestion.